# Learning Mixtures of Linear Dynamical Systems (MoLDS) via Hybrid Tensor–EM Method

### v1

@@ Abstract
Linear dynamical systems (LDSs) have been powerful tools for modeling high-dimensional time-series data across many domains, including neuroscience. However, a single LDS often struggles to capture the heterogeneity of neural data, where trajectories recorded under different conditions can have variations in their dynamics. Mixtures of linear dynamical systems (MoLDS) provide a path to model these variations in temporal dynamics for different observed trajectories. However, MoLDS remains difficult to apply in complex and noisy settings, limiting its practical use in neural data analysis. Tensor-based moment methods can provide global identifiability guarantees for MoLDS, but their performance degrades under realistic noise and complexity. Commonly used expectation-maximization (EM) methods offer flexibility in fitting latent models but are highly sensitive to initialization and prone to poor local minima. Here, we propose a tensor-based moment method that provides identifiability guarantees for learning MoLDS, which can be followed by EM updates to combine the strengths of both approaches. The novelty in our approach lies in the construction of moment tensors using the input–output data, on which we then apply Simultaneous Matrix Diagonalization (SMD) to recover globally consistent estimates of mixture weights and system parameters. These estimates can then be refined through a full Kalman Filter-Smoother EM algorithm, with closed-form updates for all LDS parameters. We validate our framework on synthetic benchmarks and real-world datasets. On synthetic data, the proposed Tensor-EM method achieves more reliable recovery and improved robustness compared to either pure tensor or randomly initialized EM methods. We then analyze neural recordings from the primate somatosensory cortex while a non-human primate performs reaches in different directions. In this setting, our method successfully models and clusters different conditions as separate subsystems, which is consistent with supervised single-LDS fits for each condition. Finally, we apply this approach to another neural dataset where monkeys perform a sequential reaching task. These results demonstrate that MoLDS provides an effective framework for modeling complex neural data in different brain regions, and that Tensor-EM is a principled and reliable approach to MoLDS learning for these applications.
@@ 1 Introduction
Neuroscience experiments now produce large volumes of high-dimensional time-series datasets, calling for new computational tools to uncover the dynamical principles underlying brain function Paninski & Cunningham (2018); Stringer & Pachitariu (2024); Urai et al. (2022). These recorded data often originate from multiple, distinct underlying dynamical processes, yet the identity of the generating system at any given time is unknown. Estimating and recovering the parameters and dynamics of such latent systems from these mixed neural trajectories is a central challenge in system identification and machine learning Ljung (1998); Durbin & Koopman (2012); Bishop (2006).
@@
Classical mixture models, such as mixtures of Gaussians (MoG) Dempster et al. (1977) and mixtures of linear regressions (MLR) De Veaux (1989), provide valuable tools for modeling these heterogeneous data. However, they are primarily designed for static settings and do not explicitly capture temporal dependencies, limiting their applicability to sequential data where temporal dynamics are central. In contrast, dynamical models such as linear dynamical systems (LDS) and their extensions-(recurrent) switching LDS (SLDS) Ghahramani & Hinton (2000); Fox (2009); Linderman et al. (2017)-are suitable for time-series data modeling with latent states and regime switches. Switching models typically target a single long trajectory and are prone to solutions that contain frequent switching; they often underperform when the goal is to learn parameters of multiple distinct LDSs from independent trials, which is precisely the structure common in neural experiments.
@@
Mixtures of linear dynamical systems (MoLDS)Chen & Poor (2022); Bakshi et al. (2023) effectively address this setting by treating each trajectory as originating from a single latent LDS. Inference of MoLDS aims to uncover the number of subsystems, their parameters, and the mixture weights from collections of input–output trajectories. This formulation enables direct parameter recovery and interpretability, making it appealing for applications such as neural data analysis, where multiple distinct but repeated dynamical motifs are observed across trials.
@@
For inference of MoLDS, tensor-based moment methods are commonly employed, where high-order statistical moments of the input–output data are reorganized into structured tensors, and their decomposition provides estimates of mixture weights and LDS parameters. The appeal of these algebraic approaches lies in their global identifiability: unlike iterative optimization methods that navigate non-convex landscapes, tensor decomposition exploits the algebraic structure of moments to directly recover parameters through polynomial equation systems that admit unique solutions under ideal conditions Anandkumar et al. (2014). However, their practical performance is often limited because moment estimates become imperfect in realistic, noisy datasets, leading to degraded parameter recovery. In parallel, likelihood-based approaches such as expectation–maximization (EM) have long been widely used for fitting classical mixture models and SLDS. EM provides a flexible iterative procedure for jointly estimating latent states, mixture weights, and system parameters through local likelihood optimization. While powerful and widely adopted, EM suffers from well-known sensitivity to initialization and susceptibility to poor local minima-limitations that become particularly problematic in the MoLDS setting, where the parameter space is high-dimensional and the likelihood surface is highly multimodal.
@@
Here, we propose a hybrid Tensor-EM framework that strategically combines global initialization with local refinement for MoLDS. We first apply tensor decomposition based on Simultaneous Matrix Diagonalization (SMD) to obtain stable and accurate initial estimates. We then use these estimates to initialize a full Kalman Filter–Smoother EM procedure, which refines all parameters through closed-form updates. This hybrid approach harnesses the global identifiability of tensor methods for robust initialization, while leveraging EM’s superior local optimization, to achieve both reliability and accuracy.
@@
We validate this framework on both synthetic and real-world datasets. On synthetic benchmarks, the proposed Tensor-EM method achieves more reliable recovery and improved robustness compared to pure tensor methods and EM with random initialization. Next, we analyze neural recordings from two different experiments. (1) Recordings from monkey somatosensory cortex during center-out reaches in different directions, where Tensor-EM identifies distinct dynamical clusters corresponding to the reaching directions, matching supervised LDS fits per direction but achieved in a fully unsupervised manner. (2) Recordings from the dorsal premotor cortex while a monkey performs reaches in continuously distributed directions, where Tensor-EM succeeds in parsing the different trials into direction-specific dynamical models. These results establish MoLDS as an effective framework for modeling heterogeneous neural systems, and demonstrate that Tensor-EM provides a principled and reliable solution for learning MoLDS in both synthetic and challenging real-world settings.
@@ 2 Related Work
Mixtures models. Mixture models (e.g., MoG and MLR) capture heterogeneity but not explicit temporal structure Dempster et al. (1977); De Veaux (1989); Li & Liang (2018). MoLDS offers fundamental advantages over classical mixture models for modeling complex temporal phenomena. Importantly, MoLDS is related to MLR through lagged-input representations: by augmenting inputs with their past values, an MLR model can approximate certain temporal dependencies. Through this connection, MoLDS inherits useful algorithmic tools, including spectral tensor methods and optimization approaches Anandkumar et al. (2014); Yi et al. (2016); Pal et al. (2022); Li & Liang (2018), while maintaining its superior modeling capacity for dynamical systems.
@@
Tensor methods and EM. Tensor decomposition methods offer a principled algebraic approach to parameter estimation in latent variable models, with polynomial-time algorithms and theoretical identifiability guarantees Anandkumar et al. (2014). Recent work in tensor-based MoLDS learning has explored different moment construction strategies and decomposition algorithms. Early approaches applied Jennrich’s algorithm directly to input-output moments Bakshi et al. (2023), while more recent work incorporates temporal lag structure using the MLR reformulation and the robust tensor power method Rui & Dahleh (2025). Our approach adopts Simultaneous Matrix Diagonalization (SMD) Kuleshov et al. (2015) for tensor decomposition, which operates on whitened tensor slices and offers improved numerical stability and robustness to noise-critical advantages in the challenging MoLDS setting.
@@
The combination of tensor initialization followed by EM-style refinement has proven effective in mixture model settings. In the MLR literature, tensor methods provide globally consistent initial estimates that lie within the basin of attraction of the maximum likelihood estimator, which are then refined using alternating minimization Yi et al. (2016); Zhong et al. (2016); Chen et al. (2021). However, alternating minimization represents a simplified version of EM that uses hard assignments rather than the probabilistic responsibilities essential for handling uncertainty in noisy settings. Our work extends this paradigm to the more complex MoLDS setting by combining tensor initialization with full Kalman filter-smoother EM, including proper handling of latent state inference and closed-form parameter updates.
@@
Contributions. Our work makes both methodological and empirical contributions to MoLDS learning and shows its practical utility in neuroscience settings. Methodologically, relative to existing MoLDS tensor methods, our approach makes several key advances: (i) we employ SMD for more stable decomposition of whitened moment tensors, (ii) we provide a principled initialization strategy for noise parameters $(Q,R)$ based on residual covariances, which was missing from prior tensor-based approaches, and (iii) we integrate a complete EM procedure with responsibility-weighted sufficient statistics for all LDS parameter updates. Compared to existing tensor-alternating minimization pipelines for MLR, our method leverages the full complexity of Kalman filtering and smoothing, which is essential for addressing the temporal dependencies and uncertainty quantification required in MoLDS applications. Empirically, we demonstrate the successful applications of tensor-based MoLDS methods to complex and real-world data analysis. While prior MoLDS tensor work has been limited to synthetic evaluations, we show that our Tensor-EM framework can effectively analyze neural recordings from primate somatosensory cortex during reaching tasks, successfully identifying distinct dynamical regimes corresponding to different movement directions in a fully unsupervised manner. This represents an important step toward making MoLDS a practical tool for important applications such as neuroscience, where capturing heterogeneous dynamics across experimental conditions is a central challenge. Together, these methodological and empirical advances demonstrate that our Tensor-EM MoLDS framework provides improved robustness and accuracy in both controlled and challenging real-world settings.
@@ 3.1 Mixture of Linear Dynamical Systems (MoLDS)
In the MoLDS setting, we observe $N$ input–output trajectories $\{(u_{i,0:T_{i}-1},\,y_{i,0:T_{i}-1})\}_{i=1}^{N}$, each generated by one of $K$ latent LDS components. Let $z_{i}\in[K]$ denote the (unknown) component for trajectory $i$, drawn i.i.d. as $z_{i}\sim\mathrm{Multinomial}(p_{1:K})$, where $p_{k}\in(0,1)$ are the mixture weights with $\sum_{k=1}^{K}p_{k}=1$, indicating the probability of a trajectory being generated by component $k$. Conditional on $z_{i}=k$, the data is generated from the following LDS:
@@
with $A_{k}\!\in\!\mathbb{R}^{n\times n}$, $B_{k}\!\in\!\mathbb{R}^{n\times m}$, $C_{k}\!\in\!\mathbb{R}^{p\times n}$, $D_{k}\!\in\!\mathbb{R}^{p\times m}$, and $Q_{k}\succeq 0$, $R_{k}\succeq 0$. The goal of MoLDS learning is to recover the mixture weights and LDS parameters $\{p_{k},(A_{k},B_{k},C_{k},D_{k},Q_{k},R_{k})\}_{k=1}^{K}$. These parameters are identifiable only up to two natural ambiguities: the ordering of the components (permutation) and similarity transformations of the latent state realization (which leave the input–output behavior unchanged).  [^1] Formally, any invertible matrix $M$ induces an equivalent realization via $A_{k}\mapsto M^{-1}A_{k}M$, $B_{k}\mapsto M^{-1}B_{k}$, $C_{k}\mapsto C_{k}M$, and $D_{k}$ unchanged, yielding the same input–output mapping.
To make this recovery possible, we adopt several standard conditions: (i) inputs are persistently exciting, (ii) each LDS component is controllable and observable, and (iii) the components are sufficiently separated to ensure identifiability Bakshi et al. (2023); Rui & Dahleh (2025).
@@ 3.2 Tensor-EM Approach Overview
Our approach (Algorithm 1) consists of two stages: a tensor initialization stage (see Algorithm 3 in the Appendix), which provides globally consistent estimates of the mixture weights and system parameters, and an EM refinement stage (see Algorithm 4), which further improves these estimates to achieve statistical efficiency. Between these two stages, a key step is the initialization of the noise parameters $(Q_{k},R_{k})$, since these are not identifiable from the tensor-based estimates. We address this gap by estimating them from the residual covariances computed using the tensor-based parameter estimates (detailed in Appendix C).
@@
7:Initialize noise parameters $\{(\hat{Q}_{k}^{(0)},\hat{R}_{k}^{(0)})\}$ from residual covariances (Appendix  C)

### v2

# Learning Mixtures of Linear Dynamical Systems via Hybrid Tensor-EM Method

@@ Abstract
Linear dynamical systems (LDSs) have been powerful tools for modeling high-dimensional time-series data across many domains, including neuroscience. However, a single LDS often struggles to capture the heterogeneity of neural data, where trajectories recorded under different conditions can have variations in their dynamics. Mixtures of linear dynamical systems (MoLDS) provide a path to model these variations in temporal dynamics for different observed trajectories. However, MoLDS remains difficult to apply in complex and noisy settings, limiting its practical use in neural data analysis. Tensor-based moment methods can provide global identifiability guarantees for MoLDS, but their practical performance degrades in high-noise or complex scenarios. Commonly used expectation-maximization (EM) methods offer flexibility in fitting latent models but are highly sensitive to initialization and prone to poor local minima. Here, we propose a tensor-based moment method that provides identifiability guarantees for learning MoLDS, which can be followed by EM updates to combine the strengths of both approaches. The novelty in our approach lies in the construction of moment tensors using the input-output data, on which we then apply Simultaneous Matrix Diagonalization (SMD) to recover globally consistent estimates of mixture weights and system parameters. These estimates can then be refined through a full Kalman EM algorithm, with closed-form updates for all LDS parameters. We validate our framework on synthetic benchmarks and real-world datasets. On synthetic data, the proposed Tensor-EM method achieves more reliable recovery and improved robustness compared to either pure tensor or randomly initialized EM methods. We then apply this method to two neural datasets from non-human primates doing reaching tasks. For both datasets, our method successfully models and clusters different conditions as separate subsystems. These results demonstrate that MoLDS provides an effective framework for modeling complex neural data in different brain regions, and that Tensor-EM is a principled and reliable approach to MoLDS learning for these applications.
@@ 1 Introduction
Neuroscience experiments now produce large volumes of high-dimensional time-series datasets, calling for new computational tools to uncover the dynamical principles underlying brain function (Paninski & Cunningham, 2018; Stringer & Pachitariu, 2024; Urai et al., 2022; Saxena & Cunningham, 2019). These recorded data often originate from multiple, distinct underlying dynamical processes, yet the identity of the generating systems at any given time is unknown. Estimating and recovering the parameters and dynamics of such latent systems from these mixed trajectories is a central challenge in system identification and machine learning (Ljung, 1998; Durbin & Koopman, 2012; Bishop, 2006).
@@
Classical mixture models, such as mixtures of Gaussians (MoG) (Dempster et al., 1977) and mixtures of linear regressions (MLR) (De Veaux, 1989), provide valuable tools for modeling these heterogeneous data. However, they are primarily designed for static settings and do not explicitly capture temporal dependencies, limiting their applicability to sequential data where temporal dynamics are central. In contrast, dynamical models such as linear dynamical systems (LDS) and their modern extensions, such as the (recurrent) switching LDS (SLDS) and the decomposed LDS (dLDS) (Ghahramani & Hinton, 2000; Fox, 2009; Linderman et al., 2017; Mudrik et al., 2024; Zhang & Saxena, 2024), are suitable for time-series data modeling with latent states and potential regime switches. Switching models typically target a single long trajectory and are prone to solutions that contain frequent switching; they are not suitable when the goal is to learn parameters of multiple distinct LDSs from independent trials, which is precisely the common structure in neural experiments.
@@
Mixtures of linear dynamical systems (MoLDS) (Chen & Poor, 2022; Bakshi et al., 2023; Rui & Dahleh, 2025) effectively address this setting by treating each trajectory as originating from a single latent LDS. Then the learning of MoLDS aims to uncover the number of subsystems, their parameters, and the mixture weights from collections of input-output trajectories. This formulation enables direct parameter recovery and interpretability, making it appealing for many applications, including large-scale neural data analysis. Neural datasets usually have large numbers of recordings across behavioral or task conditions, with many trials under each condition. An important question is whether neural populations reuse common latent dynamics across conditions or exhibit distinct dynamics for different behaviors (Vyas et al., 2020; Athalye et al., 2023). MoLDS can provide a principled way to explore this question: by identifying shared and condition-specific latent dynamics, it can reveal the mixture structure of all trials in the dataset and offer a comprehensive view of the underlying dynamical structures and motifs.
@@
For the inference of MoLDS, tensor-based moment methods are commonly employed, where high-order statistical moments of the input-output data are reorganized into structured tensors, and their decomposition provides estimates of mixture weights and LDS parameters. The appeal of these algebraic approaches lies in their global identifiability: unlike iterative optimization methods that navigate non-convex landscapes, tensor decomposition exploits the algebraic structure of moments to directly recover parameters through polynomial equation systems that admit unique solutions under ideal conditions (Anandkumar et al., 2014). However, their practical performance is often limited because moment estimates become imperfect in realistic, noisy datasets, leading to degraded parameter recovery (Kuleshov et al., 2015). In parallel, likelihood-based approaches such as expectation-maximization (EM) have long been widely used for fitting classical mixture models and LDSs. In this context, EM provides a flexible iterative procedure for jointly estimating latent states, mixture weights, and system parameters through local likelihood optimization. While powerful and widely adopted, EM suffers from well-known sensitivity to initialization and susceptibility to poor local minima (Xu & Jordan, 1996; Bishop, 2006). These limitations become particularly problematic in the MoLDS setting where the parameter space is high-dimensional and the likelihood surface is highly multimodal.
@@
Here, we propose a hybrid Tensor-EM framework that strategically combines global initialization with local refinement of these two methods. We first apply tensor decomposition based on Simultaneous Matrix Diagonalization (SMD) (Kuleshov et al., 2015) to obtain stable and accurate initial estimates. We then use these estimates to initialize a full Kalman filter-smoother EM procedure, which refines all parameters through closed-form updates over all trajectories. This hybrid approach harnesses the global identifiability of tensor methods for robust initialization, while leveraging EM’s superior local optimization, to achieve both reliability and efficiency.
@@
We validate this framework on both synthetic and real-world neural datasets. On synthetic benchmarks, the proposed Tensor-EM method achieves more reliable recovery and improved robustness compared to (i) pure tensor methods and (ii) EM with random initialization. Next, we analyze neural recordings from two different experiments: (1) Recordings from monkey somatosensory cortex during center-out reaches in different directions, where Tensor-EM identifies distinct dynamical clusters corresponding to the reaching directions, matching supervised LDS fits per direction but achieved in a fully unsupervised manner; (2) Recordings from the dorsal premotor cortex while a monkey performs sequential reaches in more distributed directions, where Tensor-EM succeeds in parsing the different trials into direction-specific dynamical models. These results establish MoLDS as an effective framework for modeling heterogeneous neural systems, and demonstrate that Tensor-EM provides a principled and reliable solution for learning MoLDS in both synthetic and challenging real-world settings.
@@ 2 Related Work
Mixtures models. Classical mixture models (e.g., MoG and MLR) capture heterogeneity but not explicit temporal structure (Dempster et al., 1977; De Veaux, 1989; Li & Liang, 2018). Importantly, MoLDS is related to MLR through lagged-input representations: by augmenting inputs with their past values, an MLR model can approximate certain temporal dependencies (Rui & Dahleh, 2025). Through this connection, MoLDS inherits useful algorithmic tools, including spectral tensor methods and optimization approaches (Anandkumar et al., 2014; Yi et al., 2016; Pal et al., 2022; Li & Liang, 2018), while maintaining its superior modeling capacity for dynamical systems.
LDS models and their variants. LDS models have been widely used to model time-series data, and several extensions have been developed to better handle nonstationarity and nonlinear structures. SLDS methods (Ghahramani & Hinton, 2000; Linderman et al., 2017) model long trajectories that switch between different dynamical regimes over time, requiring joint inference of both continuous latent states and discrete mode sequences. The dLDS method (Mudrik et al., 2024; Chen et al., 2024) captures more gradual or overlapping changes by expressing dynamics as sparse, time-varying combinations of basis operators. These frameworks are designed for long, nonstationary sequences where the dynamics themselves evolve over time. In contrast, the MoLDS setting assumes that each short trajectory is well described by a single LDS drawn from a collection of LDS components. The goal then is to identify the set of latent dynamical systems and assign trials to those components, rather than modeling intra-trial dynamics changes.
@@
Tensor methods and EM. Tensor decomposition methods offer a principled algebraic approach to parameter estimation in latent variable models, with polynomial-time algorithms and theoretical identifiability guarantees (Anandkumar et al., 2014; Panagakis et al., 2021). Recent work in tensor-based MoLDS learning has explored different moment construction strategies and decomposition algorithms. Early approaches applied Jennrich’s algorithm directly to input-output moments (Bakshi et al., 2023), while more recent work incorporates temporal lag structure using the MLR reformulation and the robust tensor power method (Rui & Dahleh, 2025). Our approach adopts the Simultaneous Matrix Diagonalization (SMD) method (Kuleshov et al., 2015), which also operates on whitened tensor slices and offers improved numerical stability and robustness to noise, which are critical advantages in the challenging MoLDS setting and neural data.
@@
The combination of tensor initialization followed by EM-style refinement has been proven effective in mixture model settings. In the MLR literature, tensor methods have been shown to provide globally consistent initial estimates that lie within the basin of attraction of the maximum likelihood estimator, which are then refined using alternating minimization (Yi et al., 2016; Zhong et al., 2016; Chen et al., 2021). However, alternating minimization represents a simplified version of EM that uses hard assignments rather than the probabilistic responsibilities essential for handling uncertainty in noisy settings. Our work extends this paradigm to the more complex MoLDS setting by combining tensor initialization with more powerful Kalman filter-smoother EM, including proper handling of latent state inference and closed-form parameter updates.
@@
Contributions. Our work makes both methodological and empirical contributions to MoLDS learning and shows its practical utility in neuroscience settings. Methodologically, relative to existing MoLDS tensor methods, our approach makes several key advances: (i) we employ SMD for more stable decomposition of whitened moment tensors, (ii) we provide a principled initialization strategy for noise parameters $(Q,R)$ based on residual covariances, which were missing from prior tensor-based approaches, and (iii) we integrate a complete EM procedure with responsibility-weighted sufficient statistics for all LDS parameter updates. Compared to existing tensor-alternating minimization pipelines for MLR, our method leverages the full flexibility of Kalman filtering and smoothing, which is essential for addressing the temporal dependencies and uncertainty quantification required in real MoLDS applications.
Empirically, we demonstrate the successful applications of Tensor-EM MoLDS methods to complex and real-world data analysis. While prior MoLDS tensor work has been limited to synthetic evaluations, we show that our Tensor-EM framework can effectively analyze neural recordings from different brain regions during distinct reaching tasks and successfully identify distinct dynamical regimes corresponding to different movement directions in a fully unsupervised manner. This represents an important step toward making MoLDS a practical tool for important applications, particularly in neuroscience, where capturing heterogeneous dynamics across experimental conditions is a central challenge. Together, these methodological and empirical advances demonstrate that our Tensor-EM MoLDS framework provides improved robustness and accuracy in both controlled and challenging real-world settings.
@@ 3.1 Mixture of Linear Dynamical Systems (MoLDS)
In the MoLDS setting (Figure 1), we observe $N$ input-output trajectories $\{(u_{i,0:T_{i}-1},\,y_{i,0:T_{i}-1})\}_{i=1}^{N}$, each generated by one of $K$ latent LDS components. Let $z_{i}\in[K]$ denote the (unknown) component for trajectory $i$, drawn i.i.d. as $z_{i}\sim\mathrm{Multinomial}(p_{1:K})$, where $p_{k}\in(0,1)$ are the mixture weights with $\sum_{k=1}^{K}p_{k}=1$, indicating the probability of a trajectory being generated by component $k$. Conditional on $z_{i}=k$, the data is generated from the following LDS:
@@
with $A_{k}\!\in\!\mathbb{R}^{n\times n}$, $B_{k}\!\in\!\mathbb{R}^{n\times m}$, $C_{k}\!\in\!\mathbb{R}^{p\times n}$, $D_{k}\!\in\!\mathbb{R}^{p\times m}$, and $Q_{k}\succeq 0$, $R_{k}\succeq 0$. The goal of MoLDS learning is to recover the mixture weights and LDS parameters $\{p_{k},(A_{k},B_{k},C_{k},D_{k},Q_{k},R_{k})\}_{k=1}^{K}$. These parameters are identifiable only up to two natural ambiguities: the ordering of the components (permutation) and similarity transformations of the latent state realization (which leave the input-output behavior unchanged).  [^1] Formally, any invertible matrix $M$ inducing an equivalent realization via $A_{k}\rightarrow M^{-1}A_{k}M$, $B_{k}\rightarrow M^{-1}B_{k}$, $C_{k}\rightarrow C_{k}M$, and $D_{k}$ unchanged, will yield the same input-output mapping.
To make this recovery possible, we adopt several standard conditions: (i) inputs are persistently exciting, (ii) each LDS component is controllable and observable, and (iii) the components are sufficiently separated to ensure identifiability (Bakshi et al., 2023; Rui & Dahleh, 2025).
Figure 1: Overview of MoLDS and the application to neural data analysis.
@@ 3.2 Tensor-EM Approach Overview
Our complete approach (Algorithm 1) consists of two stages: a tensor initialization stage (see Algorithm 3 in the Appendix), which provides globally consistent estimates of the mixture weights and system parameters, and an EM refinement stage (see Algorithm 4), which further improves these estimates and achieves statistical efficiency. Between these two stages, a key step is the initialization of the noise parameters $(Q_{k},R_{k})$, since these are not identifiable from the tensor-based estimates. We address this gap by estimating them from the residual covariances computed using the tensor-based parameter estimates (detailed in Appendix D).
@@
7:Initialize noise parameters $\{(\hat{Q}_{k}^{(0)},\hat{R}_{k}^{(0)})\}$ from residual covariances (Appendix  D)